OpenStudy (anonymous):

can someone just check if i got this right please? algebra review Given that the slope of a line is -3 and the line passes though the point (-2,4), write the equation of the line.

OpenStudy (anonymous):

 my work

OpenStudy (ivettef365):

Use formula y = mx + b substitue y = 4, x = -2 and m = -3

OpenStudy (anonymous):

was i supposed to use the point-slope formula? or was i supposed to use y= mx+b ...

OpenStudy (ivettef365):

find b, then you have 4 = -3(-2)+b 4 = 6 + b 4-6 = b -2 = b

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

wait, after the thing i did? or do i erase it

OpenStudy (ivettef365):

Then you have the equation using formula y = -3x -2